The numbers have been arranged under some rule. Based on that rule which number will come on the place of Question Mark ?
1,1,2,6,24,?,720 

Grade:11

3 Answers

Karthik Kumar
7 years ago
1=1
1=1*1(multiply preceding number with 1)
2=1*2(multiply preceding number with 2)
6=2*3(multiply preceding number with 3)
24=6*4(multiply preceding number with 4)
?=24*5=(120)(multiply preceding number with 5 .then  answer is  120)
720=120*6 (cleared answer is 120
multiply preceding number with 1.2.3.4.5.6  order
NARAYANA
7 years ago
MY DEAR FRIEND YOU ASKED LOGICAL THINKING QUASTION DO THESE TYPE OF QUASTION SOME TIME IS REQUIRED AT LAST  IT’S ANSWER IS 120.THANK YOU.
KALYAN
7 years ago
answer is 120 because 24*5 =  120
the logic is 1*1
1*2
2*3 in these way the series goes on 6*4 = 24 then 24*5 = 120
so answer is 120